The probability of event A is 0.56, and the probability of event B is 0.34. If A and B are independent events, then P(A and B) =
An equation that might be used to work out the values for A and B might be P(A and B) = P(A) x P(B)P(A and B) = 0.34 x 0.56 = 0.1904
Answer:
\(P(A and B)=0.1904\)
Step-by-step explanation:
It is given that the probability of an event A is :\(P(A)=0.56\) and the probability of an event B is :\(P(B)=0.34\).
Also, it is given that A and B are the independent events, then
\(P(A and B)=P(A){\times}P(B)\)
Thus, \(P(A and B)=(0.56){\times}(0.34)\)
\(P(A and B)=0.1904\).
Therefore, If A and B are the independent events, then the value of \(P(A and B)\) is \(0.1904\).